OpenStudy (anonymous):

What is the degree of 8x^2y^2 ?

OpenStudy (anonymous):

degree is the highest power of the variable in the polynomial

OpenStudy (anonymous):

so what is the highest power here..............?

OpenStudy (anonymous):

well, I know how to find the degree but i'm getting stuck on weather to add just the two and two, or add 2 and 2 and then 1 because x=1

OpenStudy (anonymous):

there is no need to add anything.... as degree here is 2 only what do u want to know....

OpenStudy (anonymous):

your supposed to add the two exponents to find the degree

OpenStudy (anonymous):

which equals 4, but then i was wondering if i am supposed to add 1 to the four becuase x=1 and there is an x in the equation!

OpenStudy (anonymous):

i dont know who has given u that direction to know the degree to add the degrees ....

OpenStudy (raden_zaikaria):

degrees is square .....

OpenStudy (anonymous):

noo, the 2 is just the exponent not the degree. You have to find the degree by adding. And my teacher did.

OpenStudy (anonymous):

u can only write the same is this manner : \[8(xy)^2\]

OpenStudy (mathstudent55):

To get the degree of a polynomial, for each term, add all the exponents of all the variables of the term. For a plain variable such as x or y, use exponent 1, since x = x^1. The degree of the polynominal is the same as the degree of the term with the highest degree. In this case you have only one term, so add all exponents.

OpenStudy (anonymous):

adding the exponents would give me 5 then right?

OpenStudy (mathstudent55):

Degree of 5xy^2z^5 = 1 + 2 + 5 = 8 Degree of x^3y^3 = 3 + 3 = 6 Degree of 2x^2y + 3xy^2 - 6xy^5 = ? First term: 2 + 1 = 3 Second term: 1 + 2 = 3 Third term: 1 + 5 = 6 Degree of polynomial is 6

OpenStudy (mathstudent55):

In your case all I see is x^2 and y^2, so 2 + 2 = 4. Degree is 4

OpenStudy (mathstudent55):

The coefficient does not contribute at all to the degree. Only the exponents of the variables do.

OpenStudy (anonymous):

Ohh, thanks. I get it!! Awesome!!

OpenStudy (mathstudent55):

You're welcome.